# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ Abstract
Cold-start rating prediction is a fundamental problem in recommender systems that has been extensively studied. Many methods have been proposed that exploit explicit relations among existing data, such as collaborative filtering, social recommendations and heterogeneous information network, to alleviate the data insufficiency issue for cold-start users and items. However, the explicit relations constructed based on data between different roles may be unreliable and irrelevant, which limits the performance ceiling of the specific recommendation task. Motivated by this, in this paper, we propose a flexible framework dubbed heterogeneous interaction rating network (HIRE). HIRE dose not solely rely on the pre-defined interaction pattern or the manually constructed heterogeneous information network. Instead, we devise a Heterogeneous Interaction Module (HIM) to jointly model the heterogeneous interactions and directly infer the important interactions via the observed data. In the experiments, we evaluate our model under three cold-start settings on three real-world datasets. The experimental results show that HIRE outperforms other baselines by a large margin. Furthermore, we visualize the inferred interactions of HIRE to confirm the contribution of our model.
@@ I Introduction
Rating prediction is a crucial task in recommender systems for estimating a user’s preference score on an item, which has wide-ranging applications [1, 2, 3, 4, 5] in e-commerce, online education, and entertainment platforms. In cold-start scenarios, where new users/items have limited information for prediction, accurate rating prediction is essential to attract and retain users (in the case of cold-users) and promote and market items (in the case of cold-items). Due to scarce ratings and interactions, prediction in cold-start scenarios presents a significant challenge, since current recommendation models heavily rely on the observed ratings and interactions to make predictions. To improve prediction accuracy, current solutions introduce explicit additional associations and side information to enrich the features of cold users and items. These approaches can be generally categorized into three lines of research.
As the most traditional approach, collaborative filtering (CF) [6, 7, 8, 9, 10, 11, 12] assumes similar users and items share similar ratings and exploits the user-item interactions to make the rating prediction. Recently, classical CF models have been empowered by neural networks to better incorporate the side information such as user profiles and item descriptions to encode user embeddings [8]. However, CF models can only model a single type of interactions and fail to generalize well to cold users/items with scarce rating interactions.
@@
Figure 1: An illustrative example: two main obstacles of existing approaches by exploiting interactions. To predict the first girl’s preference for music, relation between the girl and a boy seems to be irrelevant. The wrong label food for music is also unreliable for rating prediction. HIRE directly models heterogeneous interactions in a data-driven way.
To address the issue of extreme cold users/items, social recommendation approaches [13, 14, 15, 16, 17, 18, 19] extend the scope of interactions and introduce social connections to aid the recommendation task. The intuition behind this approach is that users’ preferences are highly influenced by their social communities and opinion leaders in social networks. Social relationships, collected from multiple sources, can be utilized to enhance modeling of user-item interactions, thereby alleviating the cold-start problem to some extent. However, the effectiveness of such approaches is limited by the usability and reliability of social relations. Firstly, abundant social relations are not always available due to privacy issues and data integration. For example, in some e-commerce scenarios, there are no explicit social relations between customers. Secondly, for specific recommendation tasks, irrelevant social relations can introduce noise and lead to inaccurate predictions. For instance, teenagers are more likely to rate action movies highly based on their friends’ opinions rather than their teachers’, indicating that the interaction between teachers and teenagers is irrelevant when it comes to rate action movies.
Furthermore, recent research has started to involve more complex interactions by constructing heterogeneous information networks (HINs) on entities other than users and items, such as tags and points of interest (POIs). HINs are expected to include richer semantic information and further enhance the quality of representations of users and items. Graph Neural Networks [20] and meta-path embedding [21] are two commonly-used approaches to extract users’ and items’ embeddings from HINs, enabling rating predictions based on such embeddings. However, HIN-based approaches rely on pre-defined graph structures and are therefore vulnerable to the noise and mistakes introduced during the graph construction process. They also suffer from the issue of irrelevance, similar to social recommendation approaches.
@@
To summarize, the homogeneity, unreliability and irrelevancy introduced by the explicit interactions present significant obstacles for existing recommendation approaches, especially for the cold-start scenarios. Fig. 1 illustrates the observed interactions in the system and the challenge of effectively exploiting them. This motivates us to reconsider how a rating model can automatically infer heterogeneous interactions and how rating prediction can be supported in an end-to-end fashion.
@@
In this paper, we propose a deep learning framework named Heterogeneous Interaction Rating nEtwork (HIRE) for rating prediction in cold-start scenarios. Distinguished from existing approaches we discussed, HIRE does not rely on side information in the data, but models the heterogeneous interactions for rating prediction in a fully data-driven way. Specifically, the core component of our HIRE model is the Heterogeneous Interaction Module (HIM), which jointly learns heterogeneous interactions from a prediction context of the cold users/items. HIM consists of three multi-head self-attention (MHSA) layers, the key building block of foundation models that exhibits powerful modeling capability of interactions in NLP tasks. However, in contrast to the usage of textual sequence, we leverage this interaction modeling capability of MHSA to capture the correlations between the users, items and attributes in the prediction context. The inherent metric learning idea of MHSA is beneficial for cold-start scenarios where the target users/items have only a few interactions. For ratings of user-item pairs to be predicted, to construct a prediction context, we design a neighborhood-based sampling strategy to select relevant users and items from the user-item rating bipartite graph. This provides an initial prior that is informative for predicting the ratings of cold users and items. The model is trained by learning the underlying heterogeneous interactions from multiple sampled prediction contexts to predict masked ratings. Regarding model optimization, unlike the complex optimization or memorization strategies employed in meta-learning-based recommendation systems [22, 23, 24], HIRE can be optimized with simple stochastic gradient descent algorithms, which are more flexible and efficient. Our accurate prediction results demonstrate the promise of holistically modeling heterogeneous interactions using attention layers for cold-start recommendation tasks.
The contributions of this paper are summarized as follows.
@@
  • We propose a novel learning framework, Heterogeneous Interaction Rating nEtwork (HIRE), for the cold-start rating prediction in recommendation systems. Instead of explicitly exploiting external interaction information, HIRE learns the interactions from heterogeneous sources in a holistic and data-driven fashion, and can deal with 3 typical cold-start scenarios, i.e., user cold-start, item cold-start and both user and item cold-start.
@@
  • We conduct substantial experimental studies on $3$ real-world datasets in the three cold scenarios to verify the effectiveness of HIRE. Compared with 4 collaborative filtering methods, a social recommendation method, 2 HIN-based methods and 3 state-of-the-art methods for cold-start, our HIRE outperforms others with higher precision, NDCG and MAP by 0.22, 0.30 and 0.22 on average. We conduct case studies, and show that HIM provides potential interpretive ability for rating prediction.
@@ II Related Work
Social recommendation approaches. Social recommendation relies on additional relationships from social media to enhance feature interactions. TrustWalker [16] proposes a random walk model to combine the interactions in trust networks among users with item-based collaborative filtering. MCCP [12] designs the random walk process on user-item bipartite graph to simulate the preference propagation to alleviate the data sparsity problem for cold-start users. LOCABLE [17] exploits social contents from both local user friendship and global ranking of user reputation, i.e., PageRank. SoRec [18] jointly factorizes user-item rating matrix and user-user social relation matrix. GraphRec [15] is a graph neural network framework that jointly models the social influence in a user-user graph and the rating opinions in a user-item graph. To better learn user and item embeddings for recommendation, DiffNet [14] simulates the social influence propagation in a social network by an influence diffusion neural network model. Leveraging social information alleviates data scarcity and sparsity in cold-start scenarios. However, these approaches rely on explicit and external social information, which can be incomplete and inaccurate.
TABLE I: Frequently-used Notations
Notation Description
$\mathcal{U}$$\mathcal{I}$ user/item set in the recommending system
$\mathcal{R}$ observed rating in the system
$u$/$i$ a user / an item
$m$/$n$ number of users/items in each context
$\bm{e}_{u}$/$\bm{e}_{i}$ one-hot embedding of the attributes of $u$/$i$
$h$ number of attributes of user and item
$f$ embedding dimension of each attribute
$e$ embedding dimension for a pair of user and item
$f_{U}^{k}$/$f_{I}^{k}$ linear transformation for the $k$-th attribute of user/item
$f_{R}$ rating linear transformation
$\bm{H}$ initial embedding of a prediction context
$\hat{R}$ predicted rating matrix
@@
Meta-learning for cold-start recommendation. Meta-learning [34, 35, 36, 37] aims to learn a meta model that can rapidly adapt to new tasks with a few training samples. Recently, meta-learning is used for cold-start recommendation where predication for one user is treated as a task. MeLU [23] adopts model-agnostic meta learning (MAML) [37] to learn a personalized model for each user given her preference on a few items. MAMO [24] also adopts MAML to learn a global parameter initialization for all users. In addition, a feature-specific memory module and a task-specific memory module are used to guide parameter personalization and fast prediction of user preference. TaNP [22] learns a Neural Process [38], an encoder-decoder model architecture for meta-learning, where the decoder is equipped with a task-adaptive mechanism for task-specific adaptation. These proposed meta model only explicitly model the interaction of user-item pairs, where the potential interactions among users are derived from model adaptation. Our approach can be generalized to a meta model when a prediction context is regarded as a task. The difference is that in predication contexts, we explicitly model extra interactions among multiple users and items.
@@ III Preliminary
Figure 2: Three cold-start scenarios for rating prediction: Entities in dotted boxes are the cold-start entities.
@@ III-A Problem Statement
A recommendation system has a set of users ${\mathcal{U}}=\{u_{1},\cdots,u_{M}\}$ and a set of items =${\mathcal{I}}=\{i_{1},\cdots,i_{N}\}$. The users and items are associated with categorical attributes, i.e., the user profiles and item descriptions. Here, we use $\bm{e}^{k}_{u}$ and $\bm{e}^{k}_{i}$ to denote the $k$-th categorical attribute of a user $u$ and an item $i$, respectively, which are represented by one-hot encoding. We also use $r_{ui}$ to denote the observed rating (preference) of user $u$ to item $i$. Given a set of observed ratings, ${\mathcal{R}}$, rating prediction is to predict unknown rating for pairs of users and items. Table I summarizes the frequently-used notations in this paper.
Figure 3: The architecture of HIRE. For ratings of cold user/items to be predicted, HIRE first samples a prediction context. Then, the model constructs the initial embedding for the prediction context and learns the heterogeneous interactions via $K$ HIMs. Finally, the output embedding of the last HIM is transformed into the predicted rating matrix.
Cold-start rating prediction is to predict cold-start users or/and items. There are three scenarios:
@@
  • User&Item Cold-Start: Both the user $u^{*}$ and the item $i^{*}$ are new arrival, i.e., $u^{*}\notin{\mathcal{U}}$, $i^{*}\notin{\mathcal{I}}$. They have only a few rating interactions with existing and other possible new items/users in the system.
Fig. 2 shows an example of the three cold-start scenarios in a product rating system, where entities in the dotted boxes are cold-start entities. In Fig. 2(a), it shows a User cold-start scenario to predict two new users’ ratings on three existing items, in Fig. 2(b), it shows an Item cold-start scenario to predict existing users’ preference on two cold items (e.g., movie and tennis racket), and in Fig. 2(c), it shows a User&Item cold-start scenario to predict the ratings of two new users on 2 new items (e.g., game and tennis racket).
Given the user set ${\mathcal{U}}$, item set ${\mathcal{I}}$ and the observed rating ${\mathcal{R}}$, in this paper, we build a deep learning model to support above cold-start rating prediction. It is important to note that by cold-start, the cold user and item to be predicted, as well as their associated ratings, are unavailable in the model training stage.
@@ III-B Multi-Head Self-Attention
Multi-head self-attention (MHSA) [39] is the building block of foundation models, such as BERT [40], GPT [41], Vision Transformer [42], achieving high performance in various natural language processing and computer vision tasks. Given a sequence of tokens $X=[x_{1},\cdots,x_{t}]\in\mathbb{R}^{t\times d}$ as input, self-attention models the interactions among the tokens by computing the attention weights $A\in\mathbb{R}^{t\times t}$ as Eq. (1)-(2). Here, $W_{Q}\in\mathbb{R}^{d\times d^{k}}$, $W_{K}\in\mathbb{R}^{d\times d^{k}}$ are linear weights that map $X$ to query matrix $Q\in\mathbb{R}^{t\times d^{k}}$ and key matrix $K\in\mathbb{R}^{t\times d^{k}}$, respectively, and $A_{ij}$ indicate the similarity between token $x_{i}$ and $x_{j}$.
$Q=XW_{Q},K=XW_{K}$ (1)
$A=\text{softmax}(\frac{QK^{T}}{\sqrt{d^{k}}})$ (2)
Self-attention (SA) fuses the embedding of the input tokens by taking $A$ as the aggregation weights, where $W_{V}\in\mathbb{R}^{d\times d^{v}}$ is a linear weight matrix.
$\text{SA}(X)=AXW_{V}$ (3)
@@
$\text{MHSA}(X)=[\text{SA}^{(1)}(X)\|\cdots\|\text{SA}^{(l)}(X)]W_{O}$ (4)
@@
$\Pi_{[1:t]}\circ\text{MHSA}(X)=\text{MHSA}(\Pi_{[1:t]}\circ X)$ (5)

### v3

@@ Abstract
Cold-start rating prediction is a fundamental problem in recommender systems that has been extensively studied. Many methods have been proposed that exploit explicit relations among existing data, such as collaborative filtering, social recommendations and heterogeneous information network, to alleviate the data insufficiency issue for cold-start users and items. However, the explicit relations constructed based on data between different entities may be unreliable and irrelevant, which limits the performance ceiling of a specific recommendation task. Motivated by this, in this paper, we propose a flexible framework dubbed heterogeneous interaction rating network (HIRE). HIRE does not solely rely on pre-defined interaction patterns or a manually constructed heterogeneous information network. Instead, we devise a Heterogeneous Interaction Module (HIM) to jointly model heterogeneous interactions and directly infer the important interactions via the observed data. In the experiments, we evaluate our framework under 3 cold-start settings on 3 real-world datasets. The experimental results show that HIRE outperforms other baselines by a large margin. Furthermore, we visualize the inferred interactions of HIRE to reveal the intuition behind our framework.
@@ I Introduction
Rating prediction is a fundamental task in recommender systems for estimating users’ preference scores on items, which has widespread applications [1, 2, 3, 4, 5] across e-commerce, online education, and entertainment platforms. In cold-start scenarios, where new users/items have limited information for prediction, accurate rating prediction is crucial for effective user retention (in the case of cold users) and item promotion (in the case of cold items). The scarcity of ratings and interactions in cold-start scenarios poses significant challenges, as current recommendation models heavily rely on observed ratings and interactions to make predictions. To improve prediction accuracy, current solutions incorporates explicit additional associations and side information to enrich the features of cold users and items. These approaches can be generally categorized into three lines of research.
Collaborative filtering (CF) [6, 7, 8, 9, 10, 11, 12], the most traditional approach, assumes similar users and items exhibit similar ratings and exploits the user-item interactions to make the rating prediction. However, CF models are limited to process single-type interactions and fail to generalize well to cold users/items with scarce rating interactions. Social recommendation approaches [13, 14, 15, 16, 17, 18, 19] extend the scope of interactions and introduce social connections into the recommendation framework. These approaches are on the basis that users’ preferences are highly influenced by their social communities and opinion leaders within social networks. However, the effectiveness of such approaches is constrained by the availability and reliability of social relations. Furthermore, recent research has explored the integration of complex interactions through Heterogeneous Information Networks (HINs), which incorporate entities beyond users and items, such as tags and points of interest (POIs). While HINs promise richer semantic information and enhance the representations of users and items, HIN-based approaches rely on pre-defined graph structures and are therefore vulnerable to the noise and mistakes introduced during graph construction. They also suffer from the issue of irrelevance, similar to social recommendation approaches. Fig. 1 illustrates the explicit interactions in the system and the challenge of effectively exploiting them.
@@
Figure 1: An illustrative example: to predict the first girl’s preference for music, relation between the girl and a boy seems to be irrelevant. The erroneous classification for cake (food) and movie (entertainment) is unreliable for rating prediction. HIRE directly models heterogeneous interactions in a data-driven way.
To address the issue of cold users/items, recent studies [20, 21, 22] employ meta learning to accommodate cold-start scenarios. MeLU [22] utilizes model-agnostic meta-learning (MAML) [23] to develop a personalized model for each user based on its preferences for a limited number of items. TaNP [20] develops a task-adaptive mechanism for learning the relevance of different users. MAMO [21] devises user preference memory to capture shared preferences. Despite these advances, these approaches capture only explicit interactions between users rating the same items, missing implicit relationships. In addition, their complicated optimization or memorization strategies in meta-learning-based recommendation will incur the barrier of implementations.
@@
To summarize, the homogeneity, unreliability and irrelevancy introduced by the explicit interactions present significant obstacles for existing recommendation approaches, especially for the cold-start scenarios. This motivates us to reconsider how a rating model can automatically infer heterogeneous interactions and how rating prediction can be supported in an end-to-end fashion.
@@
In this paper, we propose a deep learning framework named Heterogeneous Interaction Rating Network (HIRE) for rating prediction in cold-start scenarios. Distinguished from existing approaches we discussed, HIRE does not rely on side information in the data, but models the heterogeneous interactions for rating prediction in a fully data-driven way. Specifically, the core component of our HIRE model is the Heterogeneous Interaction Module (HIM), which jointly learns heterogeneous interactions from a prediction context of the cold users/items. HIM consists of three multi-head self-attention (MHSA) layers, the key building block of foundation models that exhibits powerful modeling capability of interactions in NLP tasks. However, in contrast to the usage of textual sequence, we leverage this interaction modeling capability of MHSA to capture the correlations between the users, items and attributes in the prediction context. The inherent metric learning idea of MHSA is beneficial for cold-start scenarios where the target users/items have only a few interactions. For ratings of user-item pairs to be predicted, to construct a prediction context, we design a neighborhood-based sampling strategy to select relevant users and items from the user-item rating bipartite graph. This provides an initial prior that is informative for predicting the ratings of cold users and items. The model is trained by learning the underlying heterogeneous interactions from multiple sampled prediction contexts to predict masked ratings. Our accurate prediction results demonstrate the promise of holistically modeling heterogeneous interactions using attention layers for cold-start recommendation tasks. The contributions of this paper are summarized as follows:
@@
  • We propose a novel learning framework, HIRE, for the cold-start rating prediction in recommendation systems. Instead of explicitly exploiting external interaction information, HIRE learns the interactions from heterogeneous sources in a holistic and data-driven fashion, and can deal with 3 typical cold-start scenarios, i.e., user cold-start, item cold-start and both user and item cold-start.
@@
  • We conduct substantial experimental studies on 3 real-world datasets in the 3 cold scenarios to verify the effectiveness of HIRE. Compared with 12 baselines, our HIRE outperforms others with higher Precision, NDCG and MAP by 0.21, 0.29 and 0.22 on average. We conduct case studies, and show that HIM provides potential interpretive ability for rating prediction.
@@ II Related Work
Social recommendation approaches. Social recommendation relies on additional relationships [27, 28, 29] from social media to enhance feature interactions. TrustWalker [16] proposes a random walk model to combine the interactions in trust networks among users with item-based collaborative filtering. MCCP [12] designs the random walk process on user-item bipartite graph to simulate the preference propagation to alleviate the data sparsity problem for cold-start users. LOCABLE [17] exploits social contents from both local user friendship and global ranking of user reputation, i.e., PageRank. SoRec [18] jointly factorizes user-item rating matrix and user-user social relation matrix. GraphRec [15] is a GNN framework that jointly models the social influence in a user-user graph and the rating opinions in a user-item graph. To better learn user and item embeddings for recommendation, DiffNet [14] simulates the social influence propagation in a social network by an influence diffusion neural network model. Leveraging social information alleviates data scarcity and sparsity in cold-start scenarios. However, these approaches rely on explicit and external social information, which can be incomplete and inaccurate.
@@
Cold-start recommendation. Meta-learning [37, 38, 39, 23] aims to learn a meta model that can rapidly adapt to new tasks with a few training samples. Recently, meta-learning is used for cold-start recommendation where predication for one user is treated as a task. MeLU [22] adopts model-agnostic meta learning (MAML) [23] to learn a personalized model for each user given her preference on a few items. MAMO [21] also adopts MAML to learn a global parameter initialization for all users. In addition, a feature-specific memory module and a task-specific memory module are used to guide parameter personalization and fast prediction of user preference. TaNP [20] learns a Neural Process [40], an encoder-decoder model architecture for meta-learning, where the decoder is equipped with a task-adaptive mechanism for task-specific adaptation. These proposed meta model only explicitly model the interaction of user-item pairs, where the potential interactions among users are derived from model adaptation.
Recently, Transformer-based methods [41, 42, 43] are proposed for capturing user-item interactions adaptively and globally. TransGNN [44] integrates Transformer and GNN layers to enhance the Transformer’s performance. Following the scaling law of Transformers, Large language models (LLMs) have presented strong zero-shot generalization ability in cold-start recommendation [45, 46, 47]. However, LLMs for recommendation require user and item possessing rich semantic descriptions as extra knowledge. Furthermore, the performance of LLMs degenerates in numerical prediction tasks.
@@ III Preliminary
Figure 2: 3 cold-start scenarios for rating prediction: Entities in dotted boxes are the cold-start entities.
@@ III-A Problem Statement
A recommendation system has a set of users ${\mathcal{U}}=\{u_{1},\cdots,u_{M}\}$ and a set of items =${\mathcal{I}}=\{i_{1},\cdots,i_{N}\}$. The users and items are associated with categorical attributes, i.e., the user profiles and item descriptions. Here, we use $\bm{e}^{k}_{u}$ and $\bm{e}^{k}_{i}$ to denote the $k$-th categorical attribute of a user $u$ and an item $i$, respectively, which are represented by one-hot encoding. We also use $r_{ui}$ to denote the observed rating (preference) of user $u$ to item $i$. Given a set of observed ratings, ${\mathcal{R}}$, rating prediction is to predict unknown rating for pairs of users and items.
Figure 3: The architecture of HIRE:For ratings of cold user/items to be predicted, HIRE first samples a prediction context. Then, the model constructs the initial embedding for the prediction context and learns the heterogeneous interactions via $L$ HIMs. Finally, the output embedding of the last HIM is transformed into the predicted rating matrix.
Cold-start rating prediction is to predict cold-start users or/and items. There are 3 scenarios:
@@
  • User&Item Cold-Start: Both the user $u^{*}$ and the item $i^{*}$ are new arrivals, i.e., $u^{*}\notin{\mathcal{U}}$, $i^{*}\notin{\mathcal{I}}$. They have only a few rating interactions with existing and other possible new items/users in the system.
Fig. 2 shows an example of the 3 cold-start scenarios in a product rating system, where entities in the dotted boxes are cold-start entities. In Fig. 2(a), it shows a User cold-start scenario to predict two new users’ ratings on three existing items, in Fig. 2(b), it shows an Item cold-start scenario to predict existing users’ preference on two cold items (e.g., movie and tennis racket), and in Fig. 2(c), it shows a User&Item cold-start scenario to predict the ratings of two new users on 2 new items (e.g., game and tennis racket).
Given the user set ${\mathcal{U}}$, item set ${\mathcal{I}}$ and the observed rating ${\mathcal{R}}$, in this paper, we build a deep learning framework to support above cold-start rating prediction. It is important to note that by cold-start, the cold user and item to be predicted, as well as their associated ratings, are unavailable in the model training stage.
@@ III-B Multi-Head Self-Attention
Multi-head self-attention (MHSA) [48] is the building block of foundation models, such as BERT [49], GPT [50], Vision Transformer [51], achieving high performance in various natural language processing and computer vision tasks. Given a sequence of tokens $X=[x_{1},\cdots,x_{t}]\in\mathbb{R}^{t\times d}$ as input, self-attention models the interactions among the tokens by computing the attention weights $A\in\mathbb{R}^{t\times t}$ as Eq. (1)-(2). Here, $W_{Q}\in\mathbb{R}^{d\times d^{k}}$, $W_{K}\in\mathbb{R}^{d\times d^{k}}$ are linear weights that map $X$ to query matrix $Q\in\mathbb{R}^{t\times d^{k}}$ and key matrix $K\in\mathbb{R}^{t\times d^{k}}$, respectively, $A_{ij}$ indicate the similarity between token $x_{i}$ and $x_{j}$, and $d^{k}$ and $d^{v}$ denote the dimensionality of the keys and values,
$Q=XW_{Q},K=XW_{K},$ (1)
$A=\text{softmax}(\frac{QK^{T}}{\sqrt{d^{k}}}),$ (2)
Self-attention (SA) fuses the embedding of the input tokens by taking $A$ as the aggregation weights, where $W_{V}\in\mathbb{R}^{d\times d^{v}}$ is a linear weight matrix,
$\text{SA}(X)=AXW_{V}.$ (3)
@@
$\text{MHSA}(X)=[\text{SA}^{(1)}(X)\|\cdots\|\text{SA}^{(l_{a})}(X)]W_{O}.$ (4)
@@
$\Pi_{[1:t]}\circ\text{MHSA}(X)=\text{MHSA}(\Pi_{[1:t]}\circ X).$ (5)